OpenStudy (anonymous):

3/4 times(-9/16) times 8/21

OpenStudy (anonymous):

reall need the answer

OpenStudy (anonymous):

-9/56

OpenStudy (anonymous):

\[\frac{3}{4}*\frac{-9}{16}*\frac{8}{21} = -\frac{9}{56}\]